dtermine which number in the set is a solution to the given equation 9x-8=56-7x;[0,4,10]
Answer by Simnepi(216)

to solve equations with an unknown on both sides of the equation we have to rearrange the terms, by 'shifting and balancing'(that's what algebra means!) so we have numbers on one side and unknowns on the other.
First we remove the 'x's from the side with the least 'x's on it.
since the right hand side has -7x we will 'undo' this by adding 7x.
Of course to keep things equal we have to add 7x to both sides.
So we get
9x-8+7x=56-7x+7x
Simplifying gives
16x-8=56 (because -7x+7x=0)
we now have the unknown on one side of the equation.
Now we have to move the numbers on to the other side of the equals sign.
To remove the -8 on the left hand side we have to 'undo' this by adding 8.
Again to keep things equal we have to add this to both sides of the equation.
So we get
16x-8+8=56+8
simplifying gives
16x=64
Now we divide by 16, leaving
x=4